# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 1.2 Our work
Table 1: Summary of results. Throughout this work, $n$ is the number of observations, $d$ is the number of features, $|\mathcal{A}|$ is the size of the active set, and $|\mathcal{I}|$ is the size of the inactive set. In addition, $\delta\in(0,1)$ is an upper bound on the failure probability, $\epsilon\in(0,1)$, $\mathbf{X}^{+}$ is the Moore–Penrose inverse of $\mathbf{X}$, and the parameters $\alpha,\gamma\in(0,1]$ are defined in Result 3.
@@
The complexity of our approximate quantum LARS algorithm depends on a few properties of the design matrix $\mathbf{X}$: the quantity $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$ and the parameters $\alpha$ and $\gamma$. We note that $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}\leq\|\mathbf{X}\|_{2}\|\mathbf{X}^{+}\|_{2}$ is upper bounded by the condition number of $\mathbf{X}$, which is simply a constant for well-behaved matrices. The existence of the parameter $\alpha\in(0,1]$ is often called (mutual) incoherence or strong irrepresentable condition in the literature [ZY06, HMZ08, Wai19] and measures the level of orthogonality between columns of $\mathbf{X}$. If the column space of $\mathbf{X}_{\mathcal{A}}$ is orthogonal to $\mathbf{X}_{i}$, then $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{i}\|_{1}=0$. On the other hand, if $\mathbf{X}_{i}$ is contained in the column space of $\mathbf{X}_{\mathcal{A}}$, then $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{i}\|_{1}=1$. Mutual incoherence has been considered by several previous works [DH01, EB02, Fuc04, Tro06, ZY06, MB06, Wai09]. Refs. [ZY06, MB06] provide several examples of families of matrices that satisfy mutual incoherence. Finally, the parameter $\gamma$, introduced and named by us as mutual overlap (between $\mathbf{y}$ and $\mathbf{X}$), measures the overlap between projections of the observation vector $\mathbf{y}$ and the columns of $\mathbf{X}$. We note that $\gamma\leq 1$ by Hölder’s inequality. In Section 4.4, we bound these three quantities for the case when $\mathbf{X}$ is the standard Gaussian matrix with i.i.d. $\mathcal{N}(0,1)$ entries. We show that, with high probability, $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$ is a constant (which follows easily from the condition number of Gaussian random matrices [CD05]), the mutual incoherence is at least $1/2$, and the mutual overlap is $\Omega(1/\sqrt{n\log{d}})$ (which maintains the overall complexity of the approximate algorithms).
In the case when $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$ is a constant and $\alpha$ and $\gamma$ are bounded away from 00, we retrieve the complexity $\widetilde{O}(\sqrt{n|\mathcal{I}|}/\epsilon+n|\mathcal{A}|+|\mathcal{A}|^{2})$ per iteration. For iterations when $|\mathcal{A}|=O(n)$, we obtain the overall quadratic improvement $\widetilde{O}(\sqrt{nd})$ over the classical LARS algorithm. Our approximate quantum LARS algorithm, however, depends on the design matrix being well-behaved in order to bound the quantities mentioned above. Regarding our approximate classical LARS algorithm, its complexity is very similar to the usual LARS algorithm. It is possible to obtain some advantage, though, if $\alpha,\gamma$ are bounded away from 00 and $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}=o(1)$.
Compared to the previous work of Chen and de Wolf [CdW23] on quantum algorithms for Lasso, our work is different in that we study the pathwise Lasso regression problem as the tuning parameter $\lambda$ varies, and not focus on a single point $\lambda$. Both works are thus incomparable.
@@
The rest of the paper is organized as follows. In Section 2, we introduce important notations, the computational model, and quantum subroutines that will be used in the paper. Next, in Section 3, we describe the steps to obtain a Lasso path, conditions for the uniqueness of the Lasso solution, and analyse the per-iteration time complexity of the classical LARS algorithm. In Section 4, we describe our two quantum LARS algorithms and a classical equivalent algorithm based on sampling. We discuss bounds in the noisy regime in Section 5 and summarize our work with some possible future directions in Section 6.
@@ 1.3 Related work
Seeing the importance for variable selection in the high-dimensional data context, algorithms such as best subsets, forward selection, and backward elimination are well studied and widely used to produce sparse solutions [Mao02, Mao04, WFL00, VK05, BT19, RS14, TFZB08, ZJ96, KM14, WB06]. Apart from the aforementioned feature selection algorithms, Ridge regression, proposed by Hoerl and Kennard [HK70], is another popular and well-studied method [McD09, Dor14, HK70, MS75, vW15, HKB75, SGV98, Kib03, Vin78].
In closer relation to our work, the homotopy method of Osborne et al. [OPT00b, OPT00a] and the Least Angle Regression (LARS) algorithm of Efron et al. [EHJT04] are feature selection algorithms which use a less greedy approach as compared to classical forward selection approaches. The LARS algorithm can be modified to give rise to algorithms for Lasso and for the efficient implementation of Forward Stagewise linear regression. Based on the work of Efron et al. [EHJT04], Rosset and Zhu [RZ07] gave a general characterization of loss-penalty pairs to allow for efficient generation of the full regularised coefficient paths. Mairal and Yu [MY12] later provided an upper bound on the number of linear segments of the Lasso regularisation path. They also showed that an $\epsilon$-approximate path for Lasso with at most $O(1/\sqrt{\varepsilon})$ segments can be obtained by developing an approximate homotopy algorithm based on newly defined $\epsilon$-approximate optimality conditions. Furthermore, Tibshirani [MY12] provided conditions for the uniqueness of the Lasso solution and extended the LARS algorithm to the setting where the Lasso solution is non-unique. Other works on algorithms for Lasso and its variants include Refs. [Rot04, Sto13, KKK08, GKZ18, AT16].
@@
In the quantum setting, considerable amount of work has been done on linear regression using the (ordinary/unregularised) least squares approach. Kaneko et al. [KMTY21] proposed a quantum algorithm for linear regression which outputs classical estimates of the regression coefficients. This improves upon the earlier works [WBL12, Wan17] where the regression coefficients are encoded in the amplitudes of a quantum state. Chakraborty et al. [CGJ19] devised quantum algorithms for the weighted and generalised variants of the least squares problem by using block-encoding techniques and Hamiltonian simulation [LC19] to improve the quantum liner systems solver. As an application of linear regression to machine learning tasks, Schuld et al. [SSP16] designed a quantum pattern recognition algorithm based on the ordinary least squares approach, using ideas from [HHL09] and [LMR14]. Kerenidis and Prakash [KP20] provided a quantum stochastic gradient descent algorithm for the weighted least squares problem using a quantum linear systems solver in the QRAM data structure model [Pra14, KP17], which allows quantum state preparation to be performed efficiently.
@@
While quantum linear regression has been relatively well studied in the unregularised setting, research on algorithms for regression using regularised least squares is not yet well established. In the sparse access model, Yu et al. [YGW19] presented a quantum algorithm for Ridge regression using a technique called parallel Hamiltonian simulation. This technique is then used to develop the quantum analogue of $K$-fold cross-validation [Reg70], which serves as an efficient performance estimator of Ridge regression. Other works on quantum Ridge regression algorithms in the sparse access model include [SX20, CYGL23]. Chen and de Wolf designed quantum algorithms for Lasso and Ridge regression from the perspective of empirical loss minimization. Both of their quantum algorithms output a classical vector whose loss is $\epsilon$-close to the minimum achievable loss. The authors also proved quantum query lower bounds for Lasso and Ridge regression. Around the same time, Bellante and Zanero [BZ22] gave a polynomial speedup for the classical matching-pursuit algorithm, which is a heuristic algorithm for the best subset selection model, i.e., linear regression with an $\ell_{0}$-regulariser. More recently, Chakraborty et al. [CMP23] gave the first quantum algorithms for least squares with general $\ell_{2}$-norm regularisation, which includes regularised versions of quantum ordinary least squares, quantum weighted least squares, and quantum generalised least squares. Their algorithms use block-encoding techniques and the framework of quantum singular value transformation [GSLW19], and demonstrate substantial improvement as compared to previous results on quantum Ridge regression [SX20, CYGL23, YGW19].
@@ 2.1 Notations
For a positive integer $n\in\mathbb{N}$, we use $[n]$ to denote the set $\{1,\dots,n\}$. For a vector $\mathbf{u}\in\mathbb{R}^{n}$, we denote its $i$-th entry as $u_{i}$ for $i\in[n]$. Let $\|\mathbf{u}\|_{p}\triangleq\big{(}\sum_{i=1}^{n}|x_{i}|^{p}\big{)}^{1/p}$ and by $\mathcal{D}_{\mathbf{u}}$ we denote the distribution over $[n]$ with probability density function $\mathcal{D}_{\mathbf{u}}(i)=|u_{i}|/\|\mathbf{u}\|_{1}$. Given a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$, we denote its $i$-th column as $\mathbf{X}_{i}$ for $i\in[d]$. Given $\mathcal{A}=\{i_{1},\dots,i_{k}\}\subseteq[d]$, define $\mathcal{A}^{c}=[d]\setminus\mathcal{A}$ and let $\mathbf{X}_{\mathcal{A}}\in\mathbb{R}^{n\times|\mathcal{A}|}$ be the matrix $\mathbf{X}_{\mathcal{A}}=[\mathbf{X}_{i_{1}},\dots,\mathbf{X}_{i_{k}}]$ formed by the columns of $\mathbf{X}$ in $\mathcal{A}$. The previous notation naturally extends to vectors, i.e., $\mathbf{u}_{\mathcal{A}}=[u_{i_{1}},\dots,u_{i_{k}}]$. Given $\mathcal{A}\subseteq[d]$ and $S\subseteq[n]$, $\mathbf{X}_{S\mathcal{A}}$ is the submatrix of $\mathbf{X}$ with rows in $S$ and columns in $\mathcal{A}$. We denote by $\operatorname{col}(\mathbf{X})$, $\operatorname{row}(\mathbf{X})$, $\operatorname{null}(\mathbf{X})$, and $\operatorname{rank}(\mathbf{X})$ the column space, row space, null space, and the rank, respectively, of $\mathbf{X}$. We denote by $\mathbf{X}^{+}$ the Moore-Penrose inverse of $\mathbf{X}$. If $\mathbf{X}$ has linearly independent columns, then $\mathbf{X}^{+}=(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}$. Note that $\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+}$ is the projection matrix onto $\operatorname{col}(\mathbf{X}_{\mathcal{A}})$. Let $\|\mathbf{X}\|_{p}\triangleq\sup_{\mathbf{u}\in\mathbb{R}^{d}:\|\mathbf{u}\|_{p}=1}\|\mathbf{X}\mathbf{u}\|_{p}$ be the $p$-norm of $\mathbf{X}$, $p\in[1,\infty]$. As special cases, $\|\mathbf{X}\|_{1}=\max_{j\in[d]}\sum_{i=1}^{n}|X_{ij}|$ and $\|\mathbf{X}\|_{\infty}=\max_{i\in[n]}\sum_{j=1}^{d}|X_{ij}|$. Let also $\lVert\mathbf{X}\rVert_{\max}\triangleq\max_{i\in[n],j\in[d]}|X_{ij}|$. We use $|{\bar{0}}\rangle$ to denote the state $|{0}\rangle\otimes\cdots\otimes|{0}\rangle$, where the number of qubits is clear from the context. We use $\widetilde{O}(\cdot)$ to hide polylogarithmic factors, i.e., $\widetilde{O}(f(n))=O(f(n)\cdot\operatorname{poly}\log(f(n)))$. See Appendix A for a summary of symbols.

### v2

@@ 1.2 Our work
Table 1: Summary of results. Throughout this work, $n$ is the number of observations, $d$ is the number of features, $\mathcal{A}$ is the active set, and $\mathcal{I}$ is the inactive set. In addition, $\epsilon>0$ is the approximation error, $\delta\in(0,1)$ is the failure probability, $T$ is the number of kinks in the regularisation path $\mathcal{P}$, $\mathbf{X}_{\mathcal{A}}=[\mathbf{X}_{i}]_{i\in\mathcal{A}}\in\mathbb{R}^{n\times|\mathcal{A}|}$ is the matrix formed by the columns of $\mathbf{X}$ in $\mathcal{A}$, $\mathbf{X}^{+}_{\mathcal{A}}$ is the Moore–Penrose inverse of $\mathbf{X}_{\mathcal{A}}$, and the parameters $\alpha_{\mathcal{A}},\gamma_{\mathcal{A}}\in(0,1)$ are defined in Result 3. The notation $\widetilde{O}(\cdot)$ omits $\operatorname{poly}\log$ terms in $n$, $d$, $T$, and $\delta$.
@@
The complexity of our approximate quantum LARS algorithm depends on a few properties of the design matrix $\mathbf{X}$: the quantity $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}$ and the parameters $\alpha_{\mathcal{A}}$ and $\gamma_{\mathcal{A}}$. We note that $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}\leq\|\mathbf{X}\|_{2}\|\mathbf{X}^{+}\|_{2}$ is upper bounded by the condition number of $\mathbf{X}$, which is simply a constant for well-behaved matrices. The existence of the parameter $\alpha_{\mathcal{A}}\in[0,1)$ is often called (mutual) incoherence or strong irrepresentable condition in the literature [126, 49, 114] and is related to the level of orthogonality between columns of $\mathbf{X}$: if the column space of $\mathbf{X}_{\mathcal{A}}$ is orthogonal to $\mathbf{X}_{i}$, then $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{i}\|_{1}=0$. Mutual incoherence has been considered by several previous works [26, 31, 38, 106, 126, 70, 113]. Refs. [126, 70] provide several examples of families of matrices that satisfy mutual incoherence. Finally, the parameter $\gamma_{\mathcal{A}}$, introduced and named by us as mutual overlap (between $\mathbf{y}$ and $\mathbf{X}$), measures the overlap between projections of the observation vector $\mathbf{y}$ and the columns of $\mathbf{X}$ (note that $\gamma_{\mathcal{A}}\leq 1$ by Hölder’s inequality).
[…]
Result 4 (Informal version of Lemma 31, Lemma 32, Lemma 33).
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a random matrix with rows sampled i.i.d. from the Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$ with positive-definite covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{d\times d}$. Let $\mathcal{A}\subseteq[d]$ of size $|\mathcal{A}|=O(n/\log{d})$. If $\mathbf{\Sigma}$ is “well behaved” (which includes $\mathbf{\Sigma}=\mathbf{I}$), then, with high probability,
$\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}=O\bigg{(}\sqrt{\frac{\log{d}}{n}}\bigg{)},\qquad\alpha_{\mathcal{A}}=1-\Omega(1),\qquad\text{and}\qquad\gamma_{\mathcal{A}}=\Omega\bigg{(}\frac{\|\mathbf{y}\|_{2}/\|\mathbf{y}\|_{1}}{\sqrt{\log{d}}}\bigg{)}.$
As a consequence, the approximate quantum and classical LARS algorithms output an approximate regularisation path with error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ and $T=O(n/\log{d})$ kinks with probability at least $1-\delta$ in time $\widetilde{O}((\|\mathbf{y}\|_{1}/\|\mathbf{y}\|_{2})(\sqrt{d}/\epsilon))$ and $\widetilde{O}((\|\mathbf{y}\|_{1}/\|\mathbf{y}\|_{2})(d/\epsilon^{2}))$ per iteration, respectively, up to $\operatorname{poly}\log$ term in $n$, $d$, and $\delta$. If further $\|\mathbf{y}\|_{1}/\|\mathbf{y}\|_{2}=\operatorname{poly}\log{n}$, then the quantum and classical complexities per iteration are $\widetilde{O}(\sqrt{d}/\epsilon)$ and $\widetilde{O}(d/\epsilon^{2})$, respectively.
For the exact meaning of a “well-behaved” covariance matrix $\mathbf{\Sigma}$, we refer the reader to Section 4.4. On a high level, it means that quantities like the minimum singular value of $\mathbf{\Sigma}$, the norm $\|\sqrt{\mathbf{\Sigma}}\|_{1}$, and other related properties of the conditional covariance matrix of $\mathbf{X}_{\mathcal{A}}|\mathbf{X}_{\mathcal{A}^{c}}$ are constant. Mostly special, the case $\mathbf{\Sigma}=\mathbf{I}$, which corresponds to a standard Gaussian matrix with entries sampled i.i.d. from $\mathcal{N}(0,1)$, trivially fulfills all these requirements.
Result 4 implies that the complexity of our approximate LARS algorithms, both quantum and classical, can be substantially better than the standard LARS algorithm depending on the input data. In the case when $\mathbf{X}$ is a Gaussian matrix and $\mathbf{y}$ is sparse, so that $\|\mathbf{y}\|_{1}/\|\mathbf{y}\|_{2}=\operatorname{poly}\log{n}$, the approximate quantum and classical complexities per iteration are $\widetilde{O}(\sqrt{d}/\epsilon)$ and $\widetilde{O}(d/\epsilon^{2})$, respectively, which only depend polylogarithmically on the number of samples $n$ (assuming that $n|\mathcal{A}|=O(\sqrt{d})$). The quantum runtime particularly is more than quadratically better than the usual $O(nd)$ classical runtime (at the expense of approximating the solution). The condition that $\mathbf{y}$ is sparse can be artificially generated through setting to zero some of the regression coefficients of the underlying true solution $\bm{\beta}^{\ast}$ connecting $\mathbf{X}$ and $\mathbf{y}$ (see Section 5). We note that sparse inputs have also provided speedups in other settings [6, 46, 7, 82, 64, 44, 62]. Independently, we believe that this result would be of interest to the compressed sensing community. The complexities for the case of random Gaussian design matrix $\mathbf{X}$ are summarised in Table 2.
Algorithm Error Time complexity per iteration
Classical LARS 1 None $O(nd)$
Simple quantum LARS 2 None $\widetilde{O}(n\sqrt{d}+n|\mathcal{A}|)$
Approximate classical LARS 4 $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}(d/\epsilon^{2}+n|\mathcal{A}|)$
Approximate quantum LARS 3 $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}(\sqrt{d}/\epsilon+n|\mathcal{A}|)$
Table 2: Summary of results for the case when the design matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ is a random Gaussian matrix with rows sampled i.i.d. from $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$ with well-behaved covariance matrix $\mathbf{\Sigma}\succ\mathbf{0}\in\mathbb{R}^{d\times d}$ and the vector of observations $\mathbf{y}\in\mathbb{R}^{n}$ is such that $\|\mathbf{y}\|_{1}/\|\mathbf{y}\|_{2}=\operatorname{poly}\log{n}$. These results are valid for active and inactive sets $\mathcal{A}$ and $\mathcal{I}$ of size $|\mathcal{A}|=O(n/\log{d})$ and $|\mathcal{I}|=\Theta(d)$. The notation $\widetilde{O}(\cdot)$ omits $\operatorname{poly}\log$ terms in $n$, $d$, and the failure probability $\delta$.
Our approximate LARS algorithms output a regularisation path with point-wise error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ (see Definition 1). Even though this is not a purely additive error $\epsilon$, it is still better than a purely multiplicative error $\epsilon\big{(}\frac{1}{2}\|\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}\|_{2}^{2}+\lambda\|\widetilde{\bm{\beta}}\|_{1}\big{)}$. Moreover, this kind of error is standard in the literature, see Section 5 and also [9, 78, 114].
Compared to the previous work of Chen and de Wolf [23] on quantum algorithms for Lasso, our work is different in that we study the pathwise Lasso regression problem as the tuning parameter $\lambda$ varies, and not focus on a single point $\lambda$. It looks infeasible to retrieve even a small interval $[\hat{\bm{\beta}}(\lambda),\hat{\bm{\beta}}(\lambda^{\prime})]$ of Lasso solutions via their algorithms. Both works are thus incomparable in terms of results. The output solution from the quantum algorithm of Chen and de Wolf can actually be used to initialise our algorithms instead of the point $(\lambda_{0},\hat{\bm{\beta}}(\lambda_{0})=\mathbf{0})$. Regarding quantum techniques, on the other hand, both works employ similar quantum subroutines like quantum amplitude estimation and quantum minimum finding. The fact that no advanced subroutines are needed to improve the dependence on both $n$ and $d$ should be seen as a strength of our (and [23]) work. Finally, we briefly mention that the lower bound $\Omega(\sqrt{d}/\epsilon^{1.5})$ from [23] does not contradict our quantum upper bound $\widetilde{O}(\sqrt{d}/\epsilon)$ since Chen and de Wolf assume an additive error $\epsilon$.
@@
The paper is organised as follows. In Section 2, we introduce important notations, the computational model, and useful quantum subroutines. In Section 3, we describe the steps to obtain a Lasso path, conditions for the uniqueness of the Lasso solution, and analyse the per-iteration time complexity of the classical LARS algorithm. In Section 4, we describe our two quantum LARS algorithms and a classical equivalent algorithm based on sampling. We discuss bounds in the noisy regime in Section 5 and summarise our work with possible future directions in Section 6.
@@ 1.3 Related work
Seeing the importance of variable selection in the high-dimensional data context, algorithms such as best subsets, forward selection, and backward elimination are well studied and widely used to produce sparse solutions [67, 68, 119, 112, 13, 90, 103, 124, 57, 116]. Apart from the aforementioned feature selection algorithms, Ridge regression, proposed by Hoerl and Kennard [47], is another popular and well-studied method [71, 29, 47, 75, 120, 48, 95, 54, 111].
[…]
@@
In the quantum setting, considerable amount of work has been done on linear regression using the (ordinary/unregularised) least squares approach. Kaneko et al. [58] proposed a quantum algorithm for linear regression which outputs classical estimates of the regression coefficients. This improves upon the earlier works [117, 115] where the regression coefficients are encoded in the amplitudes of a quantum state. Chakraborty et al. [16] devised quantum algorithms for the weighted and generalised variants of the least squares problem by using block-encoding techniques and Hamiltonian simulation [61] to improve the quantum liner systems solver. As an application of linear regression to machine learning tasks, Schuld et al. [99] designed a quantum pattern recognition algorithm based on the ordinary least squares approach, using ideas from [46, 65]. Kerenidis and Prakash [60] provided a quantum stochastic gradient descent algorithm for the weighted least squares problem using a quantum linear systems solver in the QRAM data structure model [82, 59], which allows for efficient quantum state preparation.
@@
While quantum linear regression has been relatively well studied in the unregularised setting, research on algorithms for regression using regularised least squares is not yet well established. In the sparse access model, Yu et al. [123] presented a quantum algorithm for Ridge regression using a technique called parallel Hamiltonian simulation to develop the quantum analogue of $K$-fold cross-validation [47], which serves as an efficient estimator of Ridge regression. Other works on quantum Ridge regression algorithms in the sparse access model include [102, 96, 24]. Chen and de Wolf [23] designed quantum algorithms for Lasso and Ridge regression from the perspective of empirical loss minimisation. Both of their quantum algorithms output a classical vector whose loss is $\epsilon$-close to the minimum achievable loss. Around the same time, Bellante and Zanero [14] gave a polynomial speedup for the classical matching-pursuit algorithm, which is a heuristic algorithm for the best subset selection model. More recently, Chakraborty et al. [17] gave the first quantum algorithms for least squares with general $\ell_{2}$-norm regularisation, which includes regularised versions of quantum ordinary least squares, quantum weighted least squares, and quantum generalised least squares. Their algorithms use block-encoding techniques and the framework of quantum singular value transformation [44], and demonstrate substantial improvement compared to previous results on quantum Ridge regression [102, 24, 123].
@@ 2.1 Notations
For a positive integer $n\in\mathbb{N}$, let $[n]\triangleq\{1,\dots,n\}$. Given a vector $\mathbf{u}\in\mathbb{R}^{n}$, we denote its $i$-th entry as $u_{i}$ for $i\in[n]$. Let $\|\mathbf{u}\|_{p}\triangleq\big{(}\sum_{i=1}^{n}|x_{i}|^{p}\big{)}^{1/p}$ and by $\mathcal{D}_{\mathbf{u}}$ we denote the distribution over $[n]$ with probability density function $\mathcal{D}_{\mathbf{u}}(i)=|u_{i}|/\|\mathbf{u}\|_{1}$. Given a matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$, we denote its $i$-th column as $\mathbf{X}_{i}$ for $i\in[d]$. Let $\mathbf{I}$ be the identity matrix. Given $\mathcal{A}=\{i_{1},\dots,i_{k}\}\subseteq[d]$, define $\mathcal{A}^{c}=[d]\setminus\mathcal{A}$ and let $\mathbf{X}_{\mathcal{A}}\in\mathbb{R}^{n\times|\mathcal{A}|}$ be the matrix $\mathbf{X}_{\mathcal{A}}=[\mathbf{X}_{i_{1}},\dots,\mathbf{X}_{i_{k}}]$ formed by the columns of $\mathbf{X}$ in $\mathcal{A}$. The previous notation naturally extends to vectors, i.e., $\mathbf{u}_{\mathcal{A}}=[u_{i_{1}},\dots,u_{i_{k}}]$. Given $\mathcal{A}\subseteq[d]$ and $S\subseteq[n]$, $\mathbf{X}_{S\mathcal{A}}$ is the submatrix of $\mathbf{X}$ with rows in $S$ and columns in $\mathcal{A}$. We denote by $\operatorname{col}(\mathbf{X})$, $\operatorname{row}(\mathbf{X})$, $\operatorname{null}(\mathbf{X})$, and $\operatorname{rank}(\mathbf{X})$ the column space, row space, null space, and the rank, respectively, of $\mathbf{X}$. We denote by $\mathbf{X}^{+}$ the Moore-Penrose inverse of $\mathbf{X}$. If $\mathbf{X}$ has linearly independent columns, $\mathbf{X}^{+}=(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}$. Note that $\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+}$ is the orthogonal projector onto $\operatorname{col}(\mathbf{X}_{\mathcal{A}})$. Let $\|\mathbf{X}\|_{p}\triangleq\sup_{\mathbf{u}\in\mathbb{R}^{d}:\|\mathbf{u}\|_{p}=1}\|\mathbf{X}\mathbf{u}\|_{p}$ be the $p$-norm of $\mathbf{X}$, $p\in[1,\infty]$. As special cases, $\|\mathbf{X}\|_{1}=\max_{j\in[d]}\sum_{i=1}^{n}|X_{ij}|$ and $\|\mathbf{X}\|_{\infty}=\max_{i\in[n]}\sum_{j=1}^{d}|X_{ij}|$. Let also $\lVert\mathbf{X}\rVert_{\max}\triangleq\max_{i\in[n],j\in[d]}|X_{ij}|$. We use $|{\bar{0}}\rangle$ to denote the state $|{0}\rangle\otimes\cdots\otimes|{0}\rangle$, where the number of qubits is clear from the context. We use $\widetilde{O}(\cdot)$ to hide polylogarithmic factors, i.e., $\widetilde{O}(f(n))=O(f(n)\cdot\operatorname{poly}\log(f(n)))$. See Appendix A for a summary of symbols.